The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

         mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
 http://www.astro.uni-bonn.de/~mcnews/

No. 35 August 5, 1999

Contents

News  1
Abstracts of 11 refereed papers 2
Conference contribution abstract 9

News

Catalog of LMC and SMC Cepheids available over the Internet

The EROS 2 microlensing search team announces release of two catalogs containing 290 LMC and 590 SMC Cepheids respectively. For each of these Cepheids period, Fourier coefficients, J2000 coordinates as well as light curves and magnitudes given in the EROS 2 filter system are indicated.

The catalogs are available from the EROS 2 Internet archive:

http://www-dapnia.cea.fr/Phys/Spp/Experiences/EROS/Cepheides/catalog cep.html

Florian Bauer

Abstracts of Refereed Papers

Ultraviolet Imaging Polarimetry of the Large Magellanic Cloud. I. Observations

A.A. Cole^{1}, K.H. Nordsieck^{1,2}, S.J. Gibson^{3}, & W.M. Harris^{2}

1 Department of Astronomy, University of Wisconsin–Madison, USA 2 Space Astronomy Laboratory, University of Wisconsin–Madison, USA 3 Department of Physics & Astronomy, University of Calgary, Canada

We have used the rocket-borne Wide-Field Imaging Survey Polarimeter (WISP) to image a 1.^{∘}5×4.^{∘}8 area of the western side of the Large Magellanic Cloud (LMC) at a wavelength of λ = 2150Å and a resolution of 1′′ × 1.′5. These are the first wide-field ultraviolet polarimetric images in astronomy. We find the UV background light of the LMC to be linearly polarized at levels ranging from our sensitivity limit of 4% to as high as ≈40%. In general, the polarization in a pixel increases as the flux decreases; the weighted mean value of polarization across the WISP field is 12.6% ±2.3%. The LMC’s diffuse UV background, in uncrowded areas, rises from a minimum of 5.6 ±3.1 × 10^{−8} erg s^{−1} cm^{−2} Å^{−1} Sr^{−1} (23.6 ±0.5 mag arcsec^{−2}) to 9.3 ±1.1 × 10^{−8} erg s^{−1} cm^{−2} Å^{−1} Sr^{−1} (23.1 ±0.2 mag arcsec^{−2}) in regions near the bright associations. We use our polarization maps to investigate the geometry of the interstellar medium in the LMC, and to search for evidence of a significant contribution of scattered light from OB associations to the diffuse galactic light of the LMC. Through a statistical analysis of our polarization map, we identify 9 regions of intense UV emission which may be giving rise to scattering halos in our image. We find that starlight from the N 11 complex and the LH 15 association are the strongest contributors to the scattered light component of the LMC’s diffuse galactic light. This region of the northwestern LMC can be thought of as a kiloparsec-scale reflection nebula in which OB stars illuminate distant dust grains, which scatter the light into our sightline. In contrast, the polarization map does not support the scattering of light from the large B2 complex in the southern WISP field; this effect may be astrophysical, or it may be the result of bias in our analysis.

Accepted by: The Astronomical Journal For preprints, contact cole@astro.wisc.edu

Ultraviolet Imaging Polarimetry of the Large Magellanic Cloud. II. Models

A.A. Cole^{1}, K. Wood^{2}, & K.H. Nordsieck^{1,3}

1 Department of Astronomy, University of Wisconsin–Madison, USA 2 Smithsonian Astrophysical Observatory, USA 3 Space Astronomy Laboratory, University of Wisconsin–Madison, USA

Motivated by new sounding-rocket wide-field polarimetric images of the Large Magellanic Cloud (Cole et al. 1999a), we have used a three-dimensional Monte Carlo radiation transfer code to investigate the escape of near-ultraviolet photons from young stellar associations embedded within a disk of dusty material (i.e., a galaxy). As photons propagate through the disk, they may be scattered or absorbed by dust. Scattered photons are polarized and tracked until they escape to be observed; absorbed photons heat the dust, which radiates isotropically in the far-infrared, where the galaxy is optically thin. The code produces four output images: near-UV and far-IR flux, and near-UV images in the linear Stokes parameters Q and U. From these images we construct simulated UV polarization maps of the LMC. We use these maps to place constraints on the star+dust geometry of the LMC and the optical properties of its dust grains. By tuning the model input parameters to produce maps that match the observed polarization maps, we derive information about the inclination of the LMC disk to the plane of the sky, and about the scattering phase function g. We compute a grid of models with i = 28^{∘}, 36^{∘}, and 45^{∘}, and g = 0.64, 0.70, 0.77, 0.83, and 0.90. The model which best reproduces the observed polarization maps has i = 36^{∘}_{−5}^{+2} and g ≈ 0.7. Because of the low signal-to-noise in the data, we cannot place firm constraints on the value of g. The highly inclined models do not match the observed centro-symmetric polarization patterns around bright OB associations, or the distribution of polarization values. Our models approximately reproduce the observed ultraviolet photopolarimetry of the western side of the LMC; however, the output images depend on many input parameters and are nonunique. We discuss some of the limitations of the models and outline future steps to be taken; our models make some predictions regarding the polarization properties of diffuse light across the rest of the LMC.

Accepted by: The Astronomical Journal For preprints, contact cole@astro.wisc.edu

SN 1987A’s Circumstellar Envelope, II: Kinematics of the Three Rings and the Diffuse Nebula

Arlin P.S. Crotts^{1} and Stephen R. Heathcote^{2}

1 Columbia University, USA 2 Cerro Tololo Interamerican Observatory, Chile

We present several different measurements of the velocities of structures within the circumstellar envelope of SN 1987A, including the inner, equatorial ring (ER), outer rings (ORs), and the diffuse nebulosity at radii ¡ 5 pc, based on CTIO 4m and HST data. A comparison of STIS and WFPC2 [N II]6583 loci for the rings show that the ER is expanding in radius at 10.5+-0.3 km/s, with the northern OR expanding along the line of sight at about 26 km/s, and for the southern OR, about 23 km/s. Similar results are found with CTIO 4m data. Accounting for inclination, the best fit to all data show both ORs with an expansion from the SN of 26 km/s. The ratio of the ER to OR velocities is nearly equal to the ratio of ER to OR radii, so the rings are roughly homologous, all having kinematic ages corresponding to about 20,000 yr before the SN explosion. This makes previously reported, large compositional differences between the ER and ORs difficult to understand. Additionally, a grid of longslit 4m/echelle spectra centered on the SN shows two velocity components over a region roughly coextensive with the outer circumstellar envelope extending about 5 pc (20 arcsec) from the SN. One component is blueshifted and the other redshifted from the SN centroid by about 10 km/s each. These features may represent a bipolar flow expanding from the SN, in which the ORs are propelled 10-15 km/s faster than that of the surrounding envelope into which they propagate. The kinematic timescale for the entire nebula is at least about 350,000 yr. The kinematics of these various structures constrain possible models for the evolution of the progenitor and its formation of a mass loss nebula.

Accepted by: The Astrophysical Journal

For preprints, contact arlin@astro.columbia.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9907367

The Optical Gravitational Lensing Experiment. Cepheids in the Magellanic Clouds.

III. Period-Luminosity-Color and Period-Luminosity Relations of Classical Cepheids

A. Udalski^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1}, I. Soszyński^{1}, P. Woźniak^{2}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present Period-Luminosity-Color and Period-Luminosity relations of classical Cepheids constructed for about 1240 Cepheids from the LMC and 2140 from the SMC. High quality BVI observations (120–360 epochs in the I-band and 15–40 in the BV-bands) were collected during the OGLE-II microlensing experiment. The I-band diagrams of the LMC show very small scatter, σ = 0.074 mag, indicating that Cepheid variables can potentially be a very good standard candle.

We compare relations of fundamental mode Cepheids from the LMC and SMC and we do not find significant differences of slopes of the Period-Luminosity-Color and Period-Luminosity relations in these galaxies. For the first overtone Cepheids a small change of the slope of Period-Luminosity relation is possible.

We determine the difference of distance moduli between the SMC and LMC with Cepheid relations and compare the result with difference obtained with other standard candles: RR Lyr and red clump stars. Results are very consistent and indicate that the values of zero points of the fundamental mode Cepheid relations are similar in these galaxies. The mean difference of distance moduli between the SMC and LMC is equal to μ_{SMC} − μ_{LMC} = 0.51 ± 0.03 mag.

We calibrate the Period-Luminosity-Color and Period-Luminosity relations for classical, fundamental mode Cepheids using the observed LMC relations and adopting the short LMC distance modulus, μ_{LMC} = 18.22 ± 0.05 mag, resulting from the recent determination with eclipsing system HV2274, RR Lyr and red clump stars.

Finally, we determine a constraint on the absolute magnitude of Cepheids by comparison of their mean V-band magnitude with that of RR Lyr stars in both Magellanic Clouds. The 10-day period, fundamental mode Cepheid is on average 4.63 ± 0.05 mag brighter than RR Lyr stars of LMC metallicity which with the most likely calibration of the brightness of RR Lyr stars yields M_{V}^{C,10} = −3.92 ± 0.09 mag.

Submitted to: Acta Astronomica 49

For preprints, contact udalski@sirius.astrouw.edu.pl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9907236

Eros variable stars: A catalog of Cepheids in the central regions of the Magellanic Clouds

The EROS collaboration

1 CEA, DSM, DAPNIA, Centre d’Études de Saclay, 91191 Gif-sur-Yvette, Cedex, France 2 Laboratoire de l’Accélérateur Linéaire, IN2P3 CNRS et Université Paris-Sud, BP 34 91898 Orsay Cedex, France 3 Institut d’Astrophysique de Paris, INSU CNRS, 98 bis Boulevard Arago, 75014 Paris, France 4 Astronomical Observatory, Copenhagen University, Juliane Maries Vej 30, 2100 Copenhagen, Denmark 5 Collège de France, PCC, IN2P3 CNRS, 11 place Marcelin Berthelot, 75231 Paris Cedex, France 6 Universidad de la Serena, Facultad de Ciencias, Departamento de Fisica, Casilla 554, La Serena, Chile 7 Department of Astronomy, Ohio State University, Columbus, OH 43210, USA 8 Physics Department, Ohio State University, Columbus, OH 43210, USA 9 Observatoire de Marseille, 2 place Le Verrier, 13248 Marseille Cedex 04, France

We present a catalog containing 290 LMC and 590 SMC Cepheids which have been obtained using the two 4k × 8k CCD cameras of the EROS 2 microlensing survey. The Cepheids were selected from 1,134,000 and 504,000 stars in the central regions of the LMC and SMC respectively, that were monitored over 150 nights between October 1996 and February 1997, at a rate of one measurement every night. For each Cepheid the light curves, period, magnitudes in the EROS 2 filter system, Fourier coefficients, J2000 coordinates and cross-identifications with objects referenced in the CDS Simbad database are presented. Finding charts of identified Cepheids in clusters NGC 1943, NGC 1958 and Bruck 56 are presented. The catalogue and the individual light–curves will be electronically available through the CDS (Strasbourg).

Submitted to: Astronomy & Astrophysics Supplement Series

For preprints, contact: florian@hep.saclay.cea.fr

Also available from the URL:

http://www-dapnia.cea.fr/Phys/Spp/Experiences/EROS/Cepheides/catalog cep.html

An Attempt to Pin Down the Instability Domain of Long-Period Variables

A. Gautschy

Astronomisches Institut der Universität Basel, Switzerland

The period-luminosity relation of Miras and semiregular variables in the Large Magellanic Cloud and in the Galaxy is used to locate their instability domain in the Hertzsprung-Russell diagram. We take advantage of the considerable width in luminosity of the relation at given period to assign masses to the observed long-period variables using stellar models on the the asymptotic giant branch and nonadiabatic pulsation computations. We study the sensitivity to chemical abundance of the position of the instability region on the Hertzsprung-Russell diagram. The mass function of the long-period variables along the AGB is discussed for Galactic and LMC variables. Finally, we contribute to the dispute on the pulsation mode of Miras and lend support to the view that, for most of the Mira variables, the pulsation in the fundamental mode is more likely.

Accepted by: Astronomy & Astrophysics, Main Journal

Preprints available from the URL: http://xxx.lanl.gov/abs/astro-ph/9907287

OGLE observations of four binary pulsars in the SMC

M.J. Coe^{1} and J.A. Orosz^{2}

1 Dept Physics and Astronomy, Southampton University, UK 2 Dept Astronomy and Astrophysics, Penn State University, USA

This paper presents analysis and interpretation of OGLE photometric data of four X-ray binary pulsar systems in the Small Magellanic Cloud: 1WGA J0054.9-7226, RX J0050.7-7316, RX J0049.1-7250, and 1SAX J0103.2-7209. In each case, the probable optical counterpart is identified on the basis of its optical colours. In the case of RX J0050.7-7316 the regular modulation of its optical light curve appears to reveal an ellipsoidal modulation with a period of 1.416 days. Using reasonable masses for the neutron star and the B star, we show that the amplitude and relative depths of the minima of the I-band light curve of RX J0050.7-7316 can be matched with an ellipsoidal model where the B star nearly fills its Roche lobe. For mass ratios in the range of 0.12 to 0.20, the corresponding best-fitting inclinations are about 55 degrees or larger. The neutron star would be eclipsed by the B star at inclinations larger than ≈ 60^{∘} for this particular mass ratio range. Thus RX J0050.7-7316 is a good candidate system for further study. In particular, we would need additional photometry in several colours, and most importantly, radial velocity data for the B star before we could draw more quantitative conclusions about the component masses.

Accepted by: Monthly Notices of the Royal Astronomical Society

For preprints, contact mjc@astro.soton.ac.uk

The Stellar Mass Spectrum in the Young Populous Cluster NGC 1866

S. van den Bergh

1 Dominion Astrophysical Observatory, Herzberg Institute of Astrophysics, National Research Council of Canada, 5071 West Saanich Road, Victoria, B.C., V8X 4M6, Canada

The young populous cluster NGC 1866 in the Large Magellanic Cloud (LMC), which is probably one of the most massive objects formed in the LMC during the last ∼ 3 Gyr, appears to have an unexpectedly high mass-to-light ratio. From its velocity dispersion Fischer et al. (1992) find its mass to be (1.35 ± 0.25) ·10^{5} M_{⊙}. The luminosity of this cluster is M_{V} = −8.93 ± 0.13, corresponding to L_{V} = (3.2 ± 0.4) · 10^{5}L_{V,⊙}. This yields M/L_{V} = 0.42 ± 0.09 in solar units. For a cluster of age 0.1 Gyr such a relatively high mass-to-light ratio requires a mass spectrum with an exponent x = 1.72 ± 0.09; or x = 1.75 ± 0.09 if mass loss by evolving stars is taken into account.

Accepted by: Publications of the Astronomical Society of the Pacific For preprints, contact vdb@dao.nrc.ca

Also available from the URL: http://xxx.lanl.gov/abs/astro-ph/9907271

Two Groups of Nearly Coeval Star Clusters in the Small Magellanic Cloud

R. Michael Rich^{1}, Michael M. Shara^{2}, S. Michael Fall^{2}, and David R. Zurek^{2}

1 Dept of Physics & Astronomy, University of California at Los Angeles, 8979 Math-Sciences Bldg, Los Angeles, CA 90095, USA 2 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA

We report new photometry of populous intermediate-age clusters in the SMC using the Hubble Space Telescope. In contrast to the accepted picture, these clusters appear to have formed in two brief intervals, one 8 ± 2 Gyr ago, and a more recent burst 2 ± 0.5 Gyr ago. When the ridgelines of the four clusters (NGC 339, 361, 416, and Kron 3) in the 8 Gyr burst are aligned, the dispersion in turnoff luminosities is < 0.2 mag, corresponding to a maximum age spread of ±0.7 Gyr. When the ridgelines of three clusters (NGC 152, 411, and 419) in the 2 Gyr burst are aligned, the maximum dispersion of 0.2 mag in turnoff luminosity corresponds to a permitted age spread of ±0.2 Gyr. Within each group of clusters, the entire cluster loci (including red giant branches and clumps) are nearly identical, consistent with indistinguishable metallicities and ages. In contrast to the wide dispersion in ages previously reported in the literature, our sample with more precise photometry and age measurements supports a burst-punctuated rather than a continuous cluster formation history for the 2 Gyr and 8 Gyr SMC clusters.

Accepted by: The Astronomical Journal For preprints, contact zurek@stsci.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9907398

Observations and Implications of the Star Formation History of the LMC

Jon A. Holtzman^{1} John S. Gallagher, III^{2}, Andrew A. Cole^{2}, Jeremy R. Mould^{3}, Carl J. Grillmair^{4}, and the WFPC2 IDT

1 New Mexico State University, Box 30001, Dept. 4500, Las Cruces NM 88003, USA 2 Department of Astronomy, University of Wisconsin – Madison, 475 N. Charter St., Madison, WI 53706, USA 3Mount Stromlo and Siding Spring Observatories, Australian National University, Private Bag, Weston Creek Post Office, ACT 2611, Australia 4SIRTF Science Center, Caltech, MS 100-22, Pasadena, CA 91125, USA

We present derivations of star formation histories based on color-magnitude diagrams of three fields in the LMC from HST/WFPC2 observations. A significant component of stars older than 4 Gyr is required to match the observed color-magnitude diagrams. Models with a dispersion-free agemetallicity relation are unable to reproduce the width of the observed main sequence; models with a range of metallicity at a given age provide a much better fit. Such models allow us to construct complete “population boxes” for the LMC based entirely on color-magnitude diagrams; remarkably, these qualitatively reproduce the age-metallicity relation observed in LMC clusters. We discuss some of the uncertainties in deriving star formation histories. We find, independently of the models, that the LMC bar field has a larger relative component of older stars than the outer fields. The main implications suggested by this study are: 1) the star formation history of field stars appears to differ from the age distribution of clusters, 2) there is no obvious evidence for bursty star formation, but our ability to measure bursts shorter in duration than ∼ 25% of any given age is limited by the statistics of the observed number of stars, 3) there may be some correlation of the star formation rate with the last close passage of the LMC/SMC/Milky Way, but there is no dramatic effect, and 4) the derived star formation history is probably consistent with observed abundances, based on recent chemical evolution models.

Accepted by: The Astronomical Journal For preprints, contact holtz@nmsu.edu

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9907259

Combined Analysis of the Binary-Lens Caustic-Crossing Event MACHO 98-SMC-1

C. Afonso et al., C. Alcock et al., S.H. Rhie et al., A. Udalski et al., M. Albrow et al. (The EROS, MACHO/GMAN, MPS, OGLE, and PLANET Collaborations, 94 authors)

We fit the data for the binary-lens microlensing event MACHO 98-SMC-1 from 5 different microlensing collaborations and find two distinct solutions characterized by binary separation d and mass ratio q: (d, q) = (0.54, 0.50) and (d, q) = (3.65, 0.36), where d is in units of the Einstein radius. However, the relative proper motion of the lens is very similar in the two solutions, 1.30 km/s/kpc and 1.48 km/s/kpc, thus confirming that the lens is in the Small Magellanic Cloud. The close binary can be either rotating or approximately static but the wide binary must be rotating at close its maximum allowed rate to be consistent with all the data. We measure limb-darkening coefficients for five bands ranging from I to V. As expected, these progressively decrease with rising wavelength. This is the first measurement of limb darkening for a metal-poor A star.

Submitted to: The Astrophysical Journal

For preprints, contact gould@astronomy.ohio-state.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9907247

Abstracts of Non-Refereed Papers

Investigating the process of star formation in young LMC star clusters

R.A.Johnson^{1}, S.F.Beaulieu^{1}, R.A.W.Elson^{1},G.Gilmore^{1},N.Tanvir^{2},B.Santiago^{3}

1 Institute of Astronomy, Madingley Road, Cambridge, UK 2 University of Hertfordshire, College Lane, Hatfield, UK 3 Universidade Federal do Rio Grande do Sul, Porto Alegre, Brazil

The rich star clusters in the Large Magellanic Cloud (LMC) are ideal for studying the process of star formation. Here we focus on the determination of age spreads amongst the massive stars in two young clusters, NGC1818 and NGC1805. We present colour magnitude diagrams (derived from HST data) for these clusters, and discuss the difficulties in age spread determination.

To appear in: Stellar Clusters and Associations: Convection, Rotation and Dynamos, May 1999, Sicily, ed. R. Pallavicini

Also available from the URL http://www.ast.cam.ac.uk/LMC